OpenStudy (anonymous):

Larry and Peggy are making decisions on their bank accounts. Larry wants to put the original money in an account with a higher interest rate. Peggy wants to put more money in as a principle amount because the more you start with, the more interest you will gain. Explain which method will result in more money.

OpenStudy (anonymous):

It is impossible to say without some numbers, and whether the interest is simple or compounding. With simple interest, they are proportional: I=Prt So double the interest rate will have the same effect as double the principle.

OpenStudy (anonymous):

I don't see how to do this, sorry. :/

OpenStudy (anonymous):

I think I can come up tih my own examples like if it were 2% or 5%... im pretty sure that's what its telling me that you can choose your own numbers

OpenStudy (anonymous):

@hero do you think you could help

OpenStudy (anonymous):

Larry and Peggy are making decisions on their bank accounts. Larry wants to put the original money in an account with a higher interest rate. Peggy wants to put more money in as a principle amount because the more you start with, the more interest you will gain. Explain which method will result in more money. Larry would prefer: f(0.4) = 20 + 20(0.4) Peggy would prefer: f(0.2) = 40 + 40(0.2)

OpenStudy (anonymous):

yes so they can both make the same money?

OpenStudy (anonymous):

Solve it. f(0.4) = 20 + 8 f(0.4) = 28 Now you solve the other one.

OpenStudy (anonymous):

Okay look, f(0.4) = 20 + 8 means he gets 8 dollars in interest. f(0.2) = 40 + 40(0.2) f(0.2) = 40 + 8 So they make the same amount of money.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

No problem.